# BWCache: Accelerating Video Diffusion Transformers through Block-Wise Caching

### v2

@@ Abstract
Recent advancements in Diffusion Transformers (DiTs) have established them as the state-of-the-art method for video generation. However, their inherently sequential denoising process results in inevitable latency, limiting real-world applicability. Existing acceleration methods either compromise visual quality due to architectural modifications or fail to reuse intermediate features at proper granularity. Our analysis reveals that DiT blocks are the primary contributors to inference latency. Across diffusion timesteps, the feature variations of DiT blocks exhibit a U-shaped pattern with high similarity during intermediate timesteps, which suggests substantial computational redundancy. In this paper, we propose Block-Wise Caching (BWCache), a training-free method to accelerate DiT-based video generation. BWCache dynamically caches and reuses features from DiT blocks across diffusion timesteps. Furthermore, we introduce a similarity indicator that triggers feature reuse only when the differences between block features at adjacent timesteps fall below a threshold, thereby minimizing redundant computations while maintaining visual fidelity. Extensive experiments on several video diffusion models demonstrate that BWCache achieves up to 2.24 speedup with comparable visual quality.
@@ 1 Introduction
Diffusion Models (DMs) Song and Ermon (2019); Ho et al. (2020); Dhariwal and Nichol (2021) have emerged as the predominant technology in Generative AI (GenAI) due to their ability to generate high-fidelity images and videos Rombach et al. (2022). Early DMs primarily adopted the U-Net architecture Ronneberger et al. (2015); Ramesh et al. (2022); Saharia et al. (2022); Blattmann et al. (2023). Recently, the success of Sora Brooks et al. (2024) has spurred the adoption of the Diffusion Transformer (DiT) architecture Peebles and Xie (2023); Bao et al. (2023) for video generation, achieving state-of-the-art performance. Video generation with DiT requires the sequential computation of multiple DiT blocks, specifically Spatial DiT (S-DiT) and Temporal DiT (T-DiT) blocks Ho et al. (2022). However, these blocks must be executed in sequence, which considerably slows down inference. As a result, accelerating DiT inference has become one of the most pressing challenges faced by GenAI.
@@
Traditional acceleration techniques for DMs primarily focus on architectural modifications such as distillation, pruning, and quantization Wang et al. (2024); Chu et al. (2024); Yang et al. (2024); Zhang et al. (2024a); Kang et al. (2024); Zhou et al. (2024). Although these methods reduce model complexity, they also compromise generation quality. More importantly, they require substantial computational resources and extensive training data to fine-tune, which is impractical for large pre-trained DMs. A training-free alternative is to cache the intermediate features of DMs to accelerate inference Ma et al. (2023); Zhang et al. (2024b); Zhao et al. (2024a); Kahatapitiya et al. (2024); Liu et al. (2024). By reusing these intermediate features across timesteps, these approaches eliminate redundant computations and significantly improve generation efficiency.
@@
Despite the proven effectiveness of caching in accelerating DM inference, two critical challenges remain unresolved. First, DiT is a complex structure, considering caching at a coarse granularity (e.g., timesteps) Lv et al. (2024); Liu et al. (2024) may result in the loss of essential information, while considering caching at a fine granularity (e.g., attention) Kahatapitiya et al. (2024); Zhao et al. (2024a); Yuan et al. (2024); Zhang et al. (2025) often fails to deliver significant acceleration. Thus, the first challenge is to identify which features are appropriate for caching. Second, many existing methods assume high similarity between features across adjacent timesteps, sharing features between them to speed up inference Zhao et al. (2024a); Lv et al. (2024); Wimbauer et al. (2024); Selvaraju et al. (2024); Liu et al. (2025a). In reality, however, feature similarity varies greatly depending on the generation task and specific inference process, and naive feature reuse can often degrade output details. Therefore, the second challenge is to determine when cached features should be reused or updated.
@@
In this paper, we propose Block-Wise Caching (BWCache), a training-free method to accelerate DiT-based video generation. This method can be seamlessly integrated into most DiT-based models as a plug-and-play component during the inference phase. The core idea is to cache the features from all DiT blocks at certain diffusion timesteps and reuse them across several subsequent steps. Specifically, as shown in Figure 2, DiT-based video generation typically involves providing a text prompt and initializing latent noise, which is then refined through multiple iterative denoising steps, and finally reconstructed into video frames using a VAE decoder. To determine when to reuse cached features, we introduce a similarity indicator based on the differences between block features from adjacent timesteps. Periodic recomputation is applied to mitigate potential latent drift. This approach significantly reduces inference time and resource consumption while preserving output quality. We evaluate BWCache across several video diffusion models, including Open-Sora Zheng et al. (2024), Open-Sora-Plan Lin et al. (2024), and Latte Ma et al. (2025). As shown in Figure 1, BWCache achieves superior performance at comparable computational costs to TeaCache Liu et al. (2024). The main contributions are summarized as follows.
@@ 2.2 Diffusion Model Acceleration
Despite the remarkable performance of DMs, their high inference costs limit their applicability in real-world applications. Existing acceleration techniques for DMs can generally be categorized into two groups. The first category relies on architectural changes, such as distillation simplifies networks Kang et al. (2024); Zhou et al. (2024); Salimans et al. (2024), pruning removes redundant parameters Yang et al. (2024); Zhang et al. (2024a); Castells et al. (2024), quantization reduces precision Wang et al. (2024); Chu et al. (2024); Huang et al. (2024a), and others. However, these approaches often require additional resources for fine-tuning or optimization. The second group consists of training-free methods that accelerate inference by caching intermediate features of diffusion models. DeepCache Ma et al. (2023) accelerates diffusion models by caching high-level features from sequential denoising timesteps. T-GATE Zhang et al. (2024b) enhances image generation efficiency by caching attention outputs. PAB Zhao et al. (2024a) leverages a U-shaped attention redundancy pattern to significantly accelerate the diffusion process. AdaCache Kahatapitiya et al. (2024) accelerates video generation by selectively caching computations based on video content. TeaCache Liu et al. (2024) utilizes timestep embeddings to dynamically determine which computations to cache to accelerate inference. Although these training-free methods enhance diffusion efficiency, they still face significant challenges in balancing generation quality with computational cost.
@@ 3.1 Preliminaries
DiT Block. The Spatial-Temporal DiT (ST-DiT) block Ho et al. (2022) follows the two-stage Transformer design to modulate the dynamics of diffusion features across both space and time. Each denoising step involves sequential processing through consecutive DiT blocks. For the -th DiT block () in timestep , given an input hidden state , the block first performs Adaptive Layer Normalization (AdaLN) Xu et al. (2019) followed by multi-head self-attention:
@@ 3.2 Analysis
Costs of Diffusion Models. The video generation process using DMs comprises multiple computational components, including the DiT blocks, timestep and text embeddings, VAE, and others. Figure 3(3(f)) shows the time distribution for each part of the video generation process. As illustrated, the DiT blocks require considerably more computation time than the other components. Furthermore, this proportion increases further as the video length and resolution grow, posing a substantial challenge to the efficiency of video generation.
Block Feature Variation. To better understand the internal behavior of DiT blocks during video generation, we introduce the relative L1 distance Liu et al. (2024) to measure the feature variations of each block between adjacent timesteps. For the -th block at the -th step, the block-wise relative L1 distance is calculated as follows:
A smaller value indicates minimal feature variation within the block between adjacent timesteps, while a larger value shows significant changes. In Figures 3(3(b)) and (3(b)), we present two videos generated from different prompts. The video generated with Prompt 1 depicts a relatively static scene, whereas the video generated with Prompt 2 exhibits more dynamic content (e.g., a bus moving closer from a distance). To quantify the changes in block features during the generation process for both videos, we employ heatmaps, as shown in Figures 3(3(d)) and (3(d)). In these heatmaps, the color intensity reflects the relative degree of blokc feature variation at different timesteps, with darker regions indicating greater differences. From the figures, we observe two key insights: 1) the colors are darker in the initial and final timesteps, indicating that block features change across timesteps; 2) compared to Figure 3(3(d)), the darker regions in Figure 3(3(d)) occupy a smaller proportion, suggesting that different prompts exert varying influences on the features of the blocks.
@@
Figure 3(3(f)) illustrates the aggregated relative L1 distance for two prompts. Generally, the differences in block features between adjacent diffusion timesteps exhibit a U-shaped pattern. In the early timesteps, the L1 distance is high, indicating rapid changes in noise predictions. The middle timesteps gradually stabilize, showing considerable redundancy in block calculations. While in the final timesteps, the aggregated L1 distance increases, suggesting a transition from structured noise to high-fidelity video. Notably, for simpler or more static scenes, the feature variations among blocks stay minimal, explaining why Prompt 1 exhibits a longer duration of stable timesteps compared to Prompt 2.
@@ 3.3 BWCache
where is defined in Eq.(5), and is the similarity threshold. Specifically, at each timestep, we calculate the relative L1 distance between each DiT block and its counterpart from the preceding timestep. After each timestep, we cache all block features and calculate the arithmetic mean of their relative L1 distances. If this mean is less than the indicator threshold , block computations can be skipped for subsequent timesteps starting from , allowing the reuse of cached features; otherwise, the block will be recomputed. A smaller threshold reduces cache reuse, while a larger threshold speeds up video generation, but may adversely affect video quality. The appropriate threshold should be chosen to optimize inference speed without compromising visual quality.
@@
| Model | Method | Visual Quality | Efficiency | |||||||
| VBench | LPIPS | SSIM | PSNR | FLOPs (P) | Speedup | Latency(s) | ||||
| Open-Sora | Original | 80.33% | - | - | - | 3.15 | 1 | 44.56 | ||
| -DiT | 78.21% | 0.5692 | 0.4811 | 11.91 | 3.09 | 1.03 | 43.26 | |||
| T-GATE | 77.61% | 0.3495 | 0.6760 | 15.50 | 2.75 | 1.19 | 37.45 | |||
| PAB | 78.10% | 0.2134 | 0.7798 | 22.02 | 2.55 | 1.38 | 32.38 | |||
| TeaCache | 79.16% | 0.1496 | 0.8104 | 22.39 | 2.40 | 1.50 | 29.64 | |||
| BWCache | 80.03% | 0.0879 | 0.8854 | 27.05 | 2.32 | 1.61 | 27.68 | |||
Original | 80.88% | - | - | - | 11.75 | 1 | 99.65 | ||
| -DiT | 77.55% | 0.5388 | 0.3736 | 13.85 | 11.74 | 1.01 | 98.66 | |||
| T-GATE | 80.15% | 0.3066 | 0.6219 | 18.80 | 9.88 | 1.18 | 84.45 | |||
| PAB | 80.30% | 0.2781 | 0.6627 | 19.49 | 8.69 | 1.36 | 73.41 | |||
| TeaCache | 80.32% | 0.1965 | 0.7502 | 21.50 | 3.13 | 4.41 | 22.62 | |||
| BWCache | 80.82% | 0.1001 | 0.8435 | 25.87 | 6.30 | 2.24 | 44.49 | |||
| Latte | Original | 78.95% | - | - | - | 3.36 | 1 | 26.90 | ||
| -DiT | 52.00% | 0.8513 | 0.1078 | 8.65 | 3.36 | 1.02 | 26.37 | |||
| T-GATE | 75.42% | 0.2612 | 0.6927 | 19.55 | 2.99 | 1.13 | 23.81 | |||
| PAB | 76.32% | 0.4625 | 0.5964 | 17.03 | 2.70 | 1.21 | 22.16 | |||
| TeaCache | 77.40% | 0.1969 | 0.7606 | 22.19 | 1.86 | 1.86 | 14.46 | |||
| BWCache | 78.28% | 0.1399 | 0.8181 | 26.46 | 1.77 | 1.90 | 14.16 | |||

### v3

@@ Abstract
Recent advancements in Diffusion Transformers (DiTs) have established them as the state-of-the-art method for video generation. However, their inherently sequential denoising process results in inevitable latency, limiting real-world applicability. Existing acceleration methods either compromise visual quality due to architectural modifications or fail to reuse intermediate features at proper granularity. Our analysis reveals that DiT blocks are the primary contributors to inference latency. Across diffusion timesteps, the feature variations of DiT blocks exhibit a U-shaped pattern with high similarity during intermediate timesteps, which suggests substantial computational redundancy. In this paper, we propose Block-Wise Caching (BWCache), a training-free method to accelerate DiT-based video generation. BWCache dynamically caches and reuses features from DiT blocks across diffusion timesteps. Furthermore, we introduce a similarity indicator that triggers feature reuse only when the differences between block features at adjacent timesteps fall below a threshold, thereby minimizing redundant computations while maintaining visual fidelity. Extensive experiments on several video diffusion models demonstrate that BWCache achieves up to 2.6 speedup with comparable visual quality. The code is available at https://github.com/hsc113/BWCache.
@@ 1 Introduction
Diffusion Models (DMs) Song and Ermon (2019); Ho et al. (2020); Dhariwal and Nichol (2021) have emerged as the predominant technology in Generative AI (GenAI) due to their ability to generate high-fidelity images and videos Rombach et al. (2022). Early DMs primarily adopted the U-Net architecture Ronneberger et al. (2015); Ramesh et al. (2022); Saharia et al. (2022); Blattmann et al. (2023). Recently, the success of Sora Brooks et al. (2024) has spurred the adoption of the Diffusion Transformer (DiT) architecture Peebles and Xie (2023); Bao et al. (2023) for video generation, achieving state-of-the-art performance. Video generation with DiT requires the sequential computation of multiple DiT blocks, including spatial and temporal DiT blocks Ho et al. (2022) However, these blocks must be executed in sequence, which considerably slows down inference. As a result, accelerating DiT inference has become one of the most pressing challenges faced by GenAI.
@@
Traditional acceleration techniques for DMs primarily focus on architectural modifications such as distillation, pruning, and quantization Wang et al. (2024); Chu et al. (2024); Yang et al. (2024); Zhang et al. (2024a); Kang et al. (2024); Zhou et al. (2024). Although these methods reduce model complexity, they also compromise generation quality. More importantly, they require substantial computational resources and extensive training data to fine-tune, which is impractical for large pre-trained DMs. A training-free alternative is to cache the intermediate features of DMs to accelerate inference Ma et al. (2023); Zhang et al. (2024b); Zhao et al. (2024b); Kahatapitiya et al. (2024); Liu et al. (2024). By reusing these intermediate features across timesteps, these approaches eliminate redundant computations and significantly improve generation efficiency.
Recent work has explored block-level temporal stability in DiTs. Skip-DiT Chen et al. (2025) addresses feature instability by introducing Long-Skip-Connections to stabilize DiT block features, enabling efficient caching. However, this method requires modifying the model architecture and retraining from scratch, which is impractical for large pre-trained models and limits their applicability. ProfilingDiT Ma et al. (2025b) indicates that redundancy is not uniformly distributed across timesteps. While their work does not focus on block-wise analysis in DiTs, their findings are conceptually aligned with our observation that mid-timestep regions exhibit higher redundancy. However, this method requires storing extensive intermediate features across multiple timesteps for profiling, leading to significant memory overhead.
@@
Despite the proven effectiveness of caching in accelerating DM inference, two critical challenges remain unresolved. First, DiT is a complex structure, considering caching at a coarse granularity (e.g., timesteps) Lv et al. (2025); Liu et al. (2024); Ma et al. (2025b) may result in the loss of essential information, while considering caching at a fine granularity (e.g., attention) Kahatapitiya et al. (2024); Zhao et al. (2024b); Yuan et al. (2024); Zhang et al. (2025) often fails to deliver significant acceleration. Thus, the first challenge is to identify which features are appropriate for caching. Second, many existing methods assume high similarity between features across adjacent timesteps, sharing features between them to speed up inference Zhao et al. (2024b); Lv et al. (2025); Wimbauer et al. (2024); Selvaraju et al. (2024); Liu et al. (2025c); Chen et al. (2025). In reality, however, feature similarity varies greatly depending on the generation task and specific inference process, and naive feature reuse can often degrade output details. Therefore, the second challenge is to determine when cached features should be reused or updated.
@@
In this paper, we propose Block-Wise Caching (BWCache), a training-free method to accelerate DiT-based video generation. This method can be seamlessly integrated into most DiT-based models as a plug-and-play component during the inference phase. The core idea is to cache the features from all DiT blocks at certain diffusion timesteps and reuse them across several subsequent steps. Specifically, as shown in Figure 2, DiT-based video generation typically involves providing a text prompt and initializing latent noise, which is then refined through multiple iterative denoising steps, and finally reconstructed into video frames using a VAE decoder. To determine when to reuse cached features, we introduce a similarity indicator based on the differences between block features from adjacent timesteps. Periodic recomputation is applied to mitigate potential latent drift. Our method is training-free and memory-efficient, using a lightweight similarity indicator to dynamically determine cache reuse without extensive feature storage, making it directly applicable to existing pre-trained models without architectural modifications. We evaluate BWCache across several video diffusion models, including Open-Sora Zheng et al. (2024), Open-Sora-Plan Lin et al. (2024a), Latte Ma et al. (2025a), Wan 2.1 Wan et al. (2025), and HunyuanVideo Kong et al. (2024). As shown in Figure 1, BWCache achieves superior performance at comparable computational costs to TeaCache Liu et al. (2024) and PAB Zhao et al. (2024b). The main contributions are summarized as follows.
@@ 2.2 Diffusion Model Acceleration
Despite the remarkable performance of DMs, their high inference costs limit their applicability in real-world applications. Existing acceleration techniques for DMs can generally be categorized into two groups. The first category relies on architectural changes, such as distillation simplifying networks Kang et al. (2024); Zhou et al. (2024); Salimans et al. (2024), pruning removes redundant parameters Yang et al. (2024); Zhang et al. (2024a); Castells et al. (2024), quantization reduces precision Wang et al. (2024); Chu et al. (2024); Huang et al. (2024a), and others. However, these approaches often require additional resources for fine-tuning or optimization. The second group consists of training-free methods that accelerate inference by caching intermediate features of diffusion models. DeepCache Ma et al. (2023) accelerates diffusion models by caching high-level features from sequential denoising timesteps. T-GATE Zhang et al. (2024b) enhances image generation efficiency by caching attention outputs. PAB Zhao et al. (2024b) leverages the attention redundancy to significantly accelerate the diffusion process. AdaCache Kahatapitiya et al. (2024) accelerates video generation by selectively caching computations based on video content. TeaCache Liu et al. (2024) utilizes timestep embeddings to dynamically determine which computations to cache to accelerate inference. Although these training-free methods enhance diffusion efficiency, they still face significant challenges in balancing generation quality with computational cost.
@@ 3.1 Preliminaries
DiT Block. The DiT block Ho et al. (2022); Peebles and Xie (2023) follows the two-stage Transformer design to modulate the dynamics of diffusion features across both space and time. Each denoising step involves sequential processing through consecutive DiT blocks, which is illustrated in Figure 3. For the -th DiT block () in timestep , given an input hidden state , the block first performs Adaptive Layer Normalization (AdaLN) Xu et al. (2019) followed by multi-head self-attention and cross-attention:
@@ 3.2 Analysis
Costs of Diffusion Models. The video generation process using DMs comprises multiple computational components, including the DiT blocks, timestep and text embeddings, VAE, and others. Figure 4(4(c)) shows the time distribution for each part of the video generation process. As illustrated, the DiT blocks require considerably more computation time than the other components. Furthermore, this proportion increases further as the video length and resolution grow, posing a substantial challenge to the efficiency of video generation.
Block Feature Variation. To better understand the internal behavior of DiT blocks, we introduce the relative L1 distance Liu et al. (2024) to measure the feature variations of each block between adjacent timesteps. For the -th block at the -th step, the block-wise relative L1 distance is calculated as follows:
A smaller value indicates minimal feature variation within the block between adjacent timesteps, while a larger value shows significant changes. In Figure 4(4(a)), we present two videos generated from different prompts. The video generated with Prompt 1 depicts a relatively static scene, whereas the video generated with Prompt 2 exhibits more dynamic content (e.g., a bus moving closer from a distance). We employ heatmaps to quantify the changes in block features during the generation process for both videos. In these heatmaps, the color intensity reflects the relative degree of block feature variation at different timesteps, with darker regions indicating greater differences. From the figures, we observe two key insights: 1) the colors are darker in the initial and final timesteps, indicating that block features change across timesteps; 2) compared to the figure below, the darker regions in the figure above occupy a smaller proportion, suggesting that different prompts exert varying influences on the features of the blocks.
@@
Figure 4(4(c)) illustrates the aggregated relative L1 distance for two prompts. Generally, the differences in block features between adjacent diffusion timesteps exhibit a U-shaped pattern. In the early timesteps, the L1 distance is high, indicating rapid changes in noise predictions. The middle timesteps gradually stabilize, showing considerable redundancy in block calculations. While in the final timesteps, the aggregated relative L1 distance increases, suggesting a transition from structured noise to high-fidelity video. Notably, for simpler or more static scenes, the feature variations among blocks stay minimal, explaining why Prompt 1 exhibits a longer duration of stable timesteps compared to Prompt 2. More detailed analysis of block feature behavior in Appendix A.
To further validate the consistency of this U-shaped pattern across different architectures and training methodologies, we analyze the aggregated relative L1 distance across five video diffusion models, as shown in Figure 5. The theoretical foundation for this U-shaped pattern can be understood through the lens of frequency domain analysis Qian et al. (2024). During early timesteps, low-frequency components recover, causing high feature variation (left U-shape). Middle timesteps stabilize these, minimizing changes (U-bottom). Final timesteps recover high-frequency details, increasing variation again (right U-shape). However, we observe that some models exhibit more pronounced early-stage feature oscillations compared to other models. This can be attributed to noise schedule and sampling implementation issues Lin et al. (2024b). Specifically, Open-Sora-Plan uses non-linear samplers (i.e. PNDM) that do not start from the last timestep (), creating a training-inference discrepancy where the model is forced to handle inputs that deviate from pure Gaussian noise, potentially causing initial timestep feature prediction instability and oscillations.
@@ 3.3 BWCache
where is defined in Eq.(5), and is the similarity threshold. The threshold represents an acceptable level of variation, and when features change more, fewer blocks are cached, naturally adapting to the scene dynamics. Specifically, at each timestep, we calculate the relative L1 distance between each DiT block and its counterpart from the preceding timestep. After each timestep, we cache all block features and calculate the arithmetic mean of their relative L1 distances. If this mean is less than the indicator threshold , block computations can be skipped for subsequent timesteps starting from , allowing the reuse of cached features; otherwise, the block will be recomputed. A smaller threshold reduces cache reuse, while a larger threshold speeds up video generation, but may adversely affect video quality. The appropriate threshold should be chosen to optimize inference speed without compromising visual quality.
@@
| Model | Method | Visual Quality | Efficiency | ||||
| VBench | LPIPS | SSIM | PSNR | Speedup | Latency(s) | ||
| Open-Sora | Original | 80.33% | - | - | - | 1 | 44.56 |
| -DiT | 78.21% | 0.5692 | 0.4811 | 11.91 | 1.03 | 43.26 | |
| T-GATE | 77.61% | 0.3495 | 0.6760 | 15.50 | 1.19 | 37.45 | |
| PAB | 78.10% | 0.2134 | 0.7798 | 22.02 | 1.38 | 32.38 | |
| TeaCache | 79.16% | 0.1496 | 0.8104 | 22.39 | 1.50 | 29.64 | |
| FasterCache | 79.21% | 0.1165 | 0.8435 | 23.99 | 1.35 | 32.03 | |
| BWCache | 80.03% | 0.0879 | 0.8854 | 27.05 | 1.61 | 27.68 | |
| Open-Sora- Plan | Original | 80.88% | - | - | - | 1 | 99.65 |
| -DiT | 77.55% | 0.5388 | 0.3736 | 13.85 | 1.01 | 98.66 | |
| T-GATE | 80.15% | 0.3066 | 0.6219 | 18.80 | 1.18 | 84.45 | |
| PAB | 80.30% | 0.2781 | 0.6627 | 19.49 | 1.36 | 73.41 | |
| TeaCache | 80.32% | 0.1965 | 0.7502 | 21.50 | 4.41 | 22.62 | |
| FasterCache | 80.61% | 0.1190 | 0.8103 | 24.06 | 1.51 | 65.79 | |
| BWCache | 80.82% | 0.1001 | 0.8435 | 25.87 | 2.24 | 44.49 | |
| Latte | Original | 78.95% | - | - | - | 1 | 26.90 |
| -DiT | 52.00% | 0.8513 | 0.1078 | 8.65 | 1.02 | 26.37 | |
| T-GATE | 75.42% | 0.2612 | 0.6927 | 19.55 | 1.13 | 23.81 | |
| PAB | 76.32% | 0.4625 | 0.5964 | 17.03 | 1.21 | 22.16 | |
| TeaCache | 77.40% | 0.1969 | 0.7606 | 22.19 | 1.86 | 14.46 | |
| FasterCache | 78.21% | 0.2442 | 0.7304 | 20.66 | 1.38 | 19.56 | |
| Skip-DiT | 75.76% | 0.1403 | 0.8087 | 25.50 | 1.65 | 16.28 | |
| BWCache | 78.28% | 0.1399 | 0.8181 | 26.46 | 1.90 | 14.16 | |
| Wan 2.1 | Original | 82.17% | - | - | - | 1 | 912 |
| PAB | 80.04% | 0.1623 | 0.7643 | 21.32 | 1.19 | 767 | |
| TeaCache | 81.73% | 0.2407 | 0.6593 | 18.62 | 1.41 | 644 | |
| BWCache | 81.99% | 0.0782 | 0.8539 | 25.86 | 2.00 | 457 | |
| HunyuanVideo | Original | 82.29% | - | - | - | 1 | 1122 |
| PAB | 81.73% | 0.1045 | 0.8341 | 26.69 | 1.08 | 1039 | |
| TeaCache | 82.13% | 0.1630 | 0.8052 | 24.37 | 2.27 | 493 | |
| BWCache | 82.48% | 0.0794 | 0.8903 | 29.91 | 2.60 | 433 | |